# Prospects of phase-adaptive cooling of levitated magnetic particles in a hollow-core photonic-crystal fibre

### v2

@@ Appendix C Experimental considerations
A linearly polarized laser beam at $1064$ nm with $3$ W of the total optical power is used to trap the magnetic particle inside the hollow core of the fibre. The mechanical quadratures of the particle can be detected by collecting the scattered light from the particle coming out of the fibre cladding as shown in the inset of Fig. 7 and focusing it onto a quadrature position sensor. Under ambient conditions of pressure, the particle’s motion is heavily damped due to collisions with the background gas. However, the system can be put in the underdamped oscillating regime once the fibre is evacuated using a vacuum chamber. In our experimental setup, at $1$ mbar, underdamped motion of the particle is confirmed by detecting a Lorentzian spectrum of motion along $y$-axis, as shown in Fig. 8. The measured spectrum provides a resonance peak at a frequency $\sim 2\pi\times 34.5$ kHz and damping rate $\sim$ $2\pi\times 0.3$ kHz. The noise floor in the spectrum has a value $17$ pm/$\sqrt{\mathrm{Hz}}$ and the additional spikes in the spectrum arise due to the electronic noise from the detectors and data collection.
Figure 7: Schematic of the experimental setup. HWP: half-wave plate, PBS: polarizing beam splitter. The fibre is placed inside a vacuum chamber capable of reaching ultra-high vacuum. A permanent magnet is mounted on a translational stage and can be moved back and forth radially relative to the fibre in order to adjust the amplitude of the magnetic field applied to the trapped particle. Insets: Scanning electron micrographs showing the cross-section of the HC-PCF with 44 $\mu$m core diameter, the 1 $\mu$m magnetic microparticle, and snapshot of an optically trapped magnetic particle inside the core of a HC-PCF captured with a high-speed camera. The motion sensor is a quadrant photodiode which detects the position quadrature of the levitated particle.
@@
Let us remark at the possible experimental strategy to implement such a phase adaptive feedback cooling mechanism, for the setup described in Fig. 7. The magnetic particle is trapped inside a hollow core fibre by using a linearly polarized laser beam of power $3$ W and wavelength $1064$ nm at an initial pressure of $10$ mbar. Due to collinear polarisation states of the tweezer arms, a standing wave pattern is formed inside the fibre. The starting point of the phase-adaptive feedback cooling is to detect the time varying position quadrature of the particle. This can be accomplished by imaging the particle by collecting the scattered light coming out of the fibre cladding [see inset of Fig. 7]. The collected light is then focused onto a quadrant position sensor from which a signal would be filtered out by a bandpass filter centered around $2\pi\times 34.5$ kHz, followed by a derivative circuit, a variable gain amplifier and a phase shifter. The output signal goes to the phase-modulator such as a Pockels cell. The Pockels cell will modulate the phase of the trapping beam in one of the tweezer arms in order to shift the standing wave pattern proportional to the instantaneous velocity of the levitated particle inside the HC-PCF.

### v3

@@ Appendix C Experimental considerations
A linearly polarized laser beam at $1064$ nm with $3$ W of the total optical power is used to trap the magnetic particle inside the hollow core of the fiber. The mechanical quadratures of the particle can be detected by collecting the scattered light from the particle coming out of the fiber cladding as shown in the inset of Fig. 8 and focusing it onto a quadrature position sensor. Under ambient conditions of pressure, the particle’s motion is heavily damped due to collisions with the background gas. However, the system can be put in the underdamped oscillating regime once the fiber is evacuated using a vacuum chamber. In our experimental setup, at $1$ mbar, underdamped motion of the particle is confirmed by detecting a Lorentzian spectrum of motion along $y$-axis, as shown in Fig. 9. The measured spectrum provides a resonance peak at a frequency $\sim 2\pi\times 34.5$ kHz and damping rate $\sim$ $2\pi\times 0.3$ kHz. The noise floor in the spectrum has a value $17$ pm/$\sqrt{\mathrm{Hz}}$ and the additional spikes in the spectrum arise due to the electronic noise from the detectors and data collection.
Figure 8: Schematic of the experimental setup. HWP: half-wave plate, PBS: polarizing beam splitter. The fiber is placed inside a vacuum chamber capable of reaching ultra-high vacuum. A permanent magnet is mounted on a translational stage and can be moved back and forth radially relative to the fiber in order to adjust the amplitude of the magnetic field applied to the trapped particle. Insets: Scanning electron micrographs showing the cross-section of the HC-PCF with 44 $\mu$m core diameter, the 1 $\mu$m magnetic microparticle, and snapshot of an optically trapped magnetic particle inside the core of a HC-PCF captured with a high-speed camera. The motion sensor is a quadrant photodiode which detects the position quadrature of the levitated particle.
@@
Let us remark at the possible experimental strategy to implement such a phase adaptive feedback cooling mechanism, for the setup described in Fig. 8. The magnetic particle is trapped inside a hollow core fiber by using a linearly polarized laser beam of power $3$ W and wavelength $1064$ nm at an initial pressure of $10$ mbar. Due to collinear polarisation states of the tweezer arms, a standing wave pattern is formed inside the fiber. The starting point of the phase-adaptive feedback cooling is to detect the time varying position quadrature of the particle. This can be accomplished by imaging the particle by collecting the scattered light coming out of the fiber cladding [see inset of Fig. 8]. The collected light is then focused onto a quadrant position sensor from which a signal would be filtered out by a bandpass filter centered around $2\pi\times 34.5$ kHz, followed by a derivative circuit, a variable gain amplifier and a phase shifter. The output signal goes to the phase-modulator such as a Pockels cell. The Pockels cell will modulate the phase of the trapping beam in one of the tweezer arms in order to shift the standing wave pattern proportional to the instantaneous velocity of the levitated particle inside the HC-PCF.
C.1 Mie scattering in axial feedback cooling
When the particle diameter becomes comparable to the trapping beam wavelength [56], it becomes essential to consider Mie scattering rather than relying solely on the Rayleigh approximation, which assumes a point-like dipole interaction. In the experimental setup described above, the magnetic particles have diameters of approximately 1–1.2 microns, causing them to span two adjacent antinodes of the standing wave trap (SWT). Under these conditions and depending on the refractive index of the material, the particle may experience equal attractive forces from each antinode and behave as a low-field seeker [57], with its center of mass naturally aligning to the node of the SWT. Previous studies on optical conveyor belts [58, 59, 60] have demonstrated that Mie scatterers can be effectively trapped and transported by moving the interference fringes.
The sign of the axial force originating from the SWT depends on whether the particle is a high-field or low-field seeker—a characteristic that is sensitive to the particle’s diameter and refractive index. Our analytical model is predicated on the existence of a restoring force acting on the particle’s center of mass due to the interference fringes in the axial direction, which gives rise to an axial eigenmode. Provided that the particle’s position and momentum quadratures are measurable, the center of mass of the axial eigenmode can be cooled.
As demonstrated above, the Mie scatterer in our experiment exhibits a well-defined axial eigenmode (see Fig. 9), thus confirming the feasibility of cooling its center-of-mass motion. For simplicity, the numerical model assumes a dipole-like gradient force on the particle’s center of mass. At low pressure, when the motion of the trapped particle becomes underdamped, its center of mass remains confined within a single interference fringe. This has been verified by measuring a noise-driven axial displacement of approximately 10 nm at 1 mbar pressure, compared to a fringe width of 532 nm. Consequently, by considering the center of mass as effectively residing at the node (or antinode) of the fringe, the restoring force can be modeled as being proportional to the gradient of the standing-wave intensity, thereby justifying the numerical simulation of center-of-mass cooling.
Here we present a simulation [61] of the axial optical force acting on a Mie scatterer. In our setup, a 1.2 $\mu$m magnetic particle with a refractive index of 2.11 is confined in the $y$-direction by the interference fringes of a SWT generated inside a HC-PCF using 0th-order non-diffracting Bessel beams at a wavelength of 1064 nm. Fig. 10 illustrates the restoring force on the particle’s center of mass for small displacements from the trap center.
Figure 10: Axial force acting on the centre of mass of a 1.2-micron size magnetic sphere trapped by standing wave inside the hollow-core of a photonic crystal fiber, with 1 W of trapping beam power at 1064 nm and considering small displacement from the center of the fringe. The refractive index of the material is taken to be 2.11 at 1064 nm. The horizontal axis is normalized to wavelength of the trapping beam in the air ($\lambda_{l}=1064$ nm).
@@ Appendix F Numerical procedure
Appendix G Laser frequency noise
The instantaneous frequency fluctuations of a trapping laser originate from its randomly varying phase. To model it, we express the amplitude of electric field of a single-mode electromagnetic field as [50]
$\mathcal{E}_{0}(t)=|\mathcal{E}_{0}|e^{i\phi_{l}(t)}\;,$ (56)
where $\phi_{l}(t)$ is a Gaussian random variable with zero mean, $\langle\phi_{l}(t)\rangle=0$ and variance $\langle\phi_{l}^{2}(t)\rangle=\sigma_{\phi_{l}}$. This phase noise affects the amplitude of the counter-propagating fields, so that the total electric field forming the standing-wave pattern inside the HCPCF is given by
$\mathcal{E}(y,t)$ $=2|\mathcal{E}_{0}|e^{i\phi_{l}(t)}\cos[k_{\ell}y+\phi(t)]\;.$ (57)
By averaging this field over many realizations and using the Gaussian noise property, $\langle e^{i\phi_{l}(t)}\rangle=e^{-\sigma_{\phi_{l}}/2}$, we obtain
$\langle\mathcal{E}(y,t)\rangle$ $=2|\mathcal{E}_{0}|\langle e^{i\phi_{l}(t)}\rangle\cos[k_{\ell}y+\phi(t)]=2|\mathcal{E}_{0}|e^{-\sigma_{\phi_{l}}/2}\cos[k_{\ell}y+\phi(t)]\;.$ (58)
Following the procedure described in Sec. II, we find that the laser phase noise rescales the trapping frequency from $\Omega$ to $\Omega e^{-\sigma_{\phi_{l}}/2}$. Consequently, the parameter $\beta$ is modified to $\beta e^{-\sigma_{\phi_{l}}}$, which leads to a corresponding change in the effective cooling rate from $\bar{\Gamma}$ to $\bar{\Gamma}e^{-\sigma_{\phi_{l}}}$. Within this framework, the detection noise amplitude $\eta$ and optimized gain factor $\mathcal{C}_{opt}$ are considered independent of the phase noise.